# Quantum Bayes Classifiers and Their Application in Image Classification

### v1

@@ III.2 SPODE-QBC
The construction method of SN-QBC based on the SPODE structure is as follows. There are $n$ feature nodes $x_{1}$, $x_{2}$, $\cdots$, $x_{n}$. Without loss of generality, assume that $x_{1}$ is the superparent feature as shown in Fig. 1b. The quantum circuit of SN-QBC consists of $n+1$ qubits, corresponding to the label node $y$ and $n$ feature nodes $x_{1}$ to $x_{n}$, and all nodes are initialized to |$\ket{0}$. Firstly, for the label node $y$, the class-prior probability $P(y=0)$ is counted and encoded by $R_{y}(\cdot)$, where the rotation angle is set as $\theta_{1}=f(P(y=0))$. Secondly, for the superparent $x_{1}$, one needs to count the class-conditional probabilities $P(x_{1}|y=0)$ and $P(x_{1}|y=1)$ and encode these probabilities by using $X$ and $CR_{y}(\cdot)$, with the controlled rotation angles being set as $\theta_{2}=f(P(x_{1}|y=0))$ and $\theta_{3}=f(P(x_{1}|y=1))$, respectively. For the remaining feature nodes $x_{2}$ to $x_{n}$, since each node has two parent nodes, four $C^{2}R_{y}(\cdot)$ are used to encode the corresponding class-conditional probabilities $P(x_{j}|y,x_{1})$, where the controlled rotation angles are set as $f(P(x_{j}|y,x_{1}))$ given the values of $y$ and $x_{1}$. That is, when the control bits are $yx_{1}=00,01,10,11$, the corresponding controlled rotation angles are set as $f(P(x_{j}=0|y=0,x_{1}=0))$, $f(P(x_{j}=0|y=0,x_{1}=1))$, $f(P(x_{j}=0|y=1,x_{1}=0))$, and $f(P(x_{j}=0|y=1,x_{1}=1))$, respectively.
For example, for a SPODE structured Bayesian network with 3 features shown in Fig. 1b, the quantum circuit of the SN-QBC can be seen in Fig. 3b. The output of the SPODE circuit is
@@ III.3 TAN-QBC
For a SN-QBC based on the TAN structure, one needs to obtain the TAN structure Bayesian network firstly. The TAN structure Bayes classifier is generated based on the maximum-weighted spanning tree algorithm Camerini-452 ; Zhou-423 , which includes the following steps Zhou-423 .
  1. 1.
    Calculate the conditional mutual information between any two nodes $x_{i}$ and $x_{i}$ using the following equation
@@
For a Bayesian network with a TAN structure, the quantum circuit of the corresponding SN-QBC can be constructed as follows. The label node $y$ and each feature node are initialized to |$\ket{0}$. Firstly, the class-prior probability $P(y=0)$ is encoded by $R_{y}(f(P(y=0)))$. Then, starting from the root node of the feature spanning tree, the class-conditional probabilities of each node are encoded layer by layer. For feature $x_{j}$, its conditional probabilities $P(x_{j}|y,x_{\text{parent}_{1}},\cdots,x_{\text{parent}_{n_{j}}})$ are encoded as the rotation angles of $C^{n}R_{y}$ gates, where $x_{\text{parent}_{1}}$, $\cdots$, $x_{\text{parent}_{n_{j}}}$ are $n_{j}$ parent nodes of $x_{j}$ on the upper layer.
Fig. 4a gives a simple example of a Bayesian network with a TAN structure. One can construct the corresponding QBC based on the given Bayesian network. The circuit of the SN-QBC based on the TAN structure in Fig. 4a is presented in Fig. 4b. The rotation angles of the label node $y$ and the root node $x_{1}$ of the spanning tree are set as the naïive QBCs shown in Sect. III.1. The rotation angles of $C^{2}R_{y}$ gates acting on $x_{3}$ are set as $f(P(x_{3}=0|yx_{1}=00))$, $f(P(x_{3}=0|yx_{1}=01))$, $f(P(x_{3}=0|yx_{1}=10))$, and $f(P(x_{3}=0|yx_{1}=11))$, respectively. While the rotation angles of $C^{2}R_{y}$ gates acting on $x_{2}$ are set as $f(P(x_{2}=0|yx_{3}=00))$, $f(P(x_{2}=0|yx_{3}=01))$, $f(P(x_{2}=0|yx_{3}=10))$, and $f(P(x_{2}=0|yx_{3}=11))$, respectively. The output of the TAN structure SN-QBC is
@@
(b) The circuit of the SN-QBC based on the TAN structure.
Figure 4: The quantum circuit of the SN-QBC with TAN structure.
III.4 SN-QBC Based on the Symmetric Relationship of Image Features
Both the naïve Bayes classifier and the semi-naïve Bayes classifier only consider the dependencies between a small number of feature nodes. Note that there are symmetric relationships among the sampled features in some images, such as the digits “6” and “9” shown in Fig. 5, where there exist symmetric relationships between features $x_{2}$ and $x_{5}$, and $x_{3}$ and $x_{4}$. These symmetric relationships of image features can be used to build Bayesian networks, which can then be used to establish corresponding SN-QBCs. The method for constructing a Bayesian network based on the symmetric relationships of image features is as follows.
  1. 1.
@@ III.4 SN-QBC Based on the Symmetric Relationship of Image Features
  2. 2.
    Consider the symmetric relationship of features in the sample images and add a directed edge between each pair of features that are symmetrical to each other.
@@
For the Bayesian network with symmetric relationships between features, in the case of not considering the label node $y$, the network will consist of several independent trees. In this case, one can construct the corresponding QBC by using the method for the QBC based on the TAN structure in Sect. III.3. That is, each node is initialized to |$\ket{0}$ and the class-prior probability $P(y=0)$ is encoded firstly. After that, each independent tree is considered separately. For each tree, starting from the root node, the class-conditional probabilities of each node are encoded layer by layer, respectively.
Figure 5: The Bayesian network based on symmetric relationships of features.
@@ IV Image Classification Based on QBCs
The key to the accuracy of a Bayesian classifier lies in feature selection and the structure of Bayesian networks. In this paper, we propose an image classification framework based on QBCs and local feature sampling, as illustrated in Fig. (6). Firstly, some local areas are selected in an image for feature sampling. Then, the sampled pixels are pooled and binarized to obtain local binary features. Finally, a QBC can be constructed based on a Bayesian network model with local features as nodes for image classification.
@@
IV.1 Local Features Sampling
For image classifications using a naïve QBC, if each pixel of the image is taken as a node of a Bayesian network, the network will be very complex. For example, a 28*28 image requires 784 nodes to represent the network. Although $n$ qubits can form a feature space with a dimension of $2^{n}$, the resources required for the quantum Bayesian networks are still relatively large. In addition, considering the complex relationship of interdependence among features, the complexity of Bayesian networks will continue to increase, which hinders the implementation of QBCs on current NISQ devices.
To reduce computational complexity and effectively utilize image features for classification, we propose the local feature sampling method. This method aims to obtain a small number of local key features from images. For image classification, background pixels shared by some images might not provide useful information for accurate classification by Bayes classifiers. The local feature sampling method can reduce the number of feature nodes in Bayesian networks, decrease the scale and computational complexity of QBCs, and diminish the influence of quantum noise, which is beneficial for the experimental implementation of quantum circuits.
@@ IV.1 Local Features Sampling
  1. 1.
    Local sampling and average pooling. Using a convolution operation on the sampling block and the convolution kernel, an image is locally sampled according to the specified block size (convolution kernel size). Next, the appropriate average pooling value $x^{\prime}_{i}$ of the feature node $x_{i}$ in the Bayesian network is obtained by applying average pooling YuWang-424 .
  2. 2.
    Feature binarization. In a classical Bayes classifier, assume that $P(x_{i}|c)\sim N(\mu_{c,i},\sigma_{c,i}^{2})$, where $\mu_{c,i}$ and $\sigma_{c,i}^{2}$ are the mean and the variance of the feature $x_{i}$ on the $c$-th class. One can obtain the class-prior probability by calculating the probability density function. However, it is challenging to replicate this process for a QBC. In this case, binarization is used to transform $x_{i}$ into 0 or 1 such that a single qubit can represent it.
    Here, we use the maximum likelihood estimation (MLE) Dutilleul-453 ; BoullE-425 method adopted in classical Bayes classifiers for obtaining class-conditional probabilities of continuous variables. For the binary classification case, this method runs as follows.
    Assume that $P(x_{i}|c_{0})\sim N(\mu_{c_{0},i},\sigma_{c_{0},i}^{2})$ and $P(x_{i}|c_{1})\sim N(\mu_{c_{1},i},\sigma_{c_{1},i}^{2})$. When two Gaussian functions have only one intersection, which is denoted as $x_{\text{ins}}$, as shown in Fig. (7a), the feature value is set as follows.
    $x_{i}=\begin{cases}0,&\text{if}\ x^{\prime}_{i}\leq x_{\text{ins}}\ \text{and}\ \mu_{c_{0},i}\leq\mu_{c_{1},i},\ \text{or}\ x^{\prime}_{i}>x_{\text{ins}}\ \text{and}\ \mu_{c_{0},i}>\mu_{c_{1},i};\\ 1,&\text{otherwise.}\end{cases}$ (19)
    where $x^{\prime}_{i}$ is the average pooling value generated by the the average pooling processing. If there are two intersection points $x_{\text{ins1}}$ and $x_{\text{ins2}}$ with $x_{\text{ins1}}\leq x_{\text{ins2}}$, as shown in Fig. (7b), in the cases where $\mu_{c_{0},i}\leq\mu_{c_{1},i}$, the feature value is set as
@@
    Conversely, when $\mu_{c_{0},i}>\mu_{c_{1},i}$, the feature value is set as
@@
By utilizing the aforementioned method, it is possible to obtain a limited number of local features. These local features serve as nodes of Bayesian networks and are utilized to construct QBCs.
IV.2 Image Classification Algorithm Based on Quantum Bayes Classifier
The image classification algorithm based on QBCs and local feature sampling is shown in Fig. (6). The algorithm consists of two stages, namely the image preprocessing stage, and the Bayesian networks and QBCs construction stage. In the image preprocessing stage, certain local areas are chosen for sampling. After that, the sampled features are pooled, and the Gaussian method is used for binarization. This process converts the sampled key features into either 0 or 1, allowing a feature to be represented by a single qubit in QBCs. In the second stage, a Bayesian network model is selected and the corresponding QBC is constructed with local key features as nodes. The algorithm runs as follows.
  1. 1.
@@ IV.2 Image Classification Algorithm Based on Quantum Bayes Classifier
  2. 2.
    Local sampling and average pooling are performed on sample images, and the mean value $\mu_{i}$ and variance $\sigma_{i}$ of the corresponding features are calculated. The Gaussian binarization method is executed to obtain the binarized features, as described in Sect. IV.1.
  3. 3.
    The QBC circuit is constructed based on the chosen Bayesian network model by using local binarized features as nodes. Also see Sect. III.
  4. 4.
    The class prior probability and class conditional probability required for the Bayesian network are calculated statistically. These values are then loaded into the QBC circuit as controlled angles to complete the construction of the QBC.
  5. 5.
    To predict the class of a new image, one needs to repeat the step 2 to obtain a new local key feature $X^{\prime}$, and then measure the probabilities of states |$\ket{0X^{\prime}}$ and |$\ket{1X^{\prime}}$ on the QBC circuit. The class with the highest probability will be the classification result of the image.
@@ V Simulations
In this paper, we test our QBCs using the MNIST LecunBottou-417 and Fashion-MNIST XiaoRasul-416 datasets. The MindQuantum Mindquantum-418 quantum simulation platform is used as the implementation platform for simulation of QBCs. The sampling block size is set to 7*7, and the average pooling method is applied. All images are sampled according to the sampling positions shown in Fig. (8) to obtain 9 local key features ($n=9$).
As is shown in Fig. (8), four Bayesian network models are used to construct four corresponding QBCs, i.e., the naïve QBC, the SN-QBC based on the SPODE structure (SPODE-QBC), the SN-QBC based on the TAN structure (TAN-QBC), and the SN-QBC based on the symmetric relationships of features (symmetric-QBC). For the TAN structures of the MNIST and Fashion-MNIST datasets, the spanning trees of the training data of the two datasets are calculated according to the maximum weighted spanning tree algorithm Camerini-452 ; Zhou-423 introduced in Sect. III.3.

### v2

@@ III.2 SPODE-QBC
For the SPODE Bayesian network, $n$ attribute nodes are $x_{1}$, $x_{2}$, $\cdots$, $x_{n}$. Without loss of generality, assume that $x_{1}$ is the super-parent node, as shown in Fig. 1b. The quantum circuit of the SN-QBC based on the SPODE structure (SPODE-QBC) consists of $n+1$ qubits, corresponding to $y$ and $x_{1}$ to $x_{n}$. The construction of SPODE-QBC is as follows:
  1. (a)
    All $n+1$ qubits are initialized to |$\ket{0}$.
  2. (b)
    For the label node $y$, the class-prior probability $P(y=0)$ is counted and encoded by $R_{y}(\cdot)$, where the rotation angle is set as $\theta_{1}=f(P(y=0))$.
  3. (c)
    For the super-parent $x_{1}$, one needs to count the class-conditional probabilities $P(x_{1}=0|y=0)$ and $P(x_{1}=0|y=1)$ and encode these probabilities by using $X$ and $CR_{y}(\cdot)$, with the controlled rotation angles being set as $\theta_{2}=f(P(x_{1}=0|y=1))$ and $\theta_{3}=f(P(x_{1}=0|y=0))$, respectively.
  4. (d)
    For the remaining attributes $x_{2}$ to $x_{n}$, since each node has two parent nodes, four $C^{2}R_{y}(\cdot)$ are used to encode the corresponding class-conditional probabilities $P(x_{j}=0|y,x_{1})$, where the controlled rotation angles are set as $f(P(x_{j}=0|y,x_{1}))$ given the values of $y$ and $x_{1}$. That is, when the control bits are $yx_{1}=00,01,10,11$, the corresponding controlled rotation angles are set as $f(P(x_{j}=0|y=0,x_{1}=0))$, $f(P(x_{j}=0|y=0,x_{1}=1))$, $f(P(x_{j}=0|y=1,x_{1}=0))$, and $f(P(x_{j}=0|y=1,x_{1}=1))$, respectively.
Figure 4: The quantum circuit of the SPODE-QBC with 3 attributes and $x_{1}$ as the super-parent node.
For example, for a SPODE structured Bayesian network with 3 attributes shown in Fig. 1b, the quantum circuit of the SPODE-QBC is depicted in Fig. 4. Specifically, the class-conditional probabilities of $x_{2}$, $x_{3}$, and $x_{4}$ are encoded as $\theta_{4}=f(P(x_{2}=0|yx_{1}=11))$, $\theta_{5}=f(P(x_{3}=0|yx_{1}=11))$, $\theta_{6}=f(P(x_{2}=0|yx_{1}=10))$, $\theta_{7}=f(P(x_{3}=0|yx_{1}=10))$, $\theta_{8}=f(P(x_{2}=0|yx_{1}=00))$, $\theta_{9}=f(P(x_{3}=0|yx_{1}=00))$, $\theta_{10}=f(P(x_{2}=0|yx_{1}=01))$, $\theta_{11}=f(P(x_{3}=0|yx_{1}=01))$. The output of the SPODE circuit is
@@ III.3 TAN-QBC
For a SN-QBC based on the TAN structure, one needs to obtain the TAN structure Bayesian network firstly. The TAN structure Bayes classifier is generated based on the maximum-weighted spanning tree algorithm Camerini-452 ; Zhou-423 , which includes the following steps Zhou-423 :
  1. (1)
    Calculate the conditional mutual information between any two nodes $x_{i}$ and $x_{j}$ using the following equation
@@
For a Bayesian network with a TAN structure, the quantum circuit of the TAN-QBC consists of $n+1$ qubits, corresponding to $y$ and $x_{1}$ to $x_{n}$. The construction of TAN-QBC is as follows:
  1. (a)
    All qubits are initialized to |$\ket{0}$.
  2. (b)
    The class-prior probability $P(y=0)$ is encoded by $R_{y}(f(P(y=0)))$.
  3. (c)
    Starting from the root node of the feature spanning tree, the class-conditional probabilities of each node are encoded layer by layer. For attribute $x_{j}$, it has at most two parent nodes, i.e., $y$ and $x_{\text{parent}_{j}}$, where $x_{\text{parent}_{j}}$ is the parent node of $x_{j}$ on the upper layer (note that the root node of the feature spanning tree only has one parent $y$). The class-conditional probabilities $P(x_{j}|y,x_{\text{parent}_{j}})$ are encoded into the circuit as the rotation angles of four $C^{2}R_{y}(\cdot)$ gates.
A simple example of a Bayesian network with a TAN structure is given in Fig. 5a. While the circuit of the TAN-QBC based on the TAN structure is presented in Fig. 5b. The rotation angles of the label node $y$ and the root node $x_{1}$ of the spanning tree are set as the naïive-QBCs shown in Sect. III.1. The rotation angles of $C^{2}R_{y}$ gates acting on $x_{3}$ are set as $\theta_{4}=f(P(x_{3}=0|yx_{1}=11))$, $\theta_{5}=f(P(x_{3}=0|yx_{1}=10))$, $\theta_{6}=f(P(x_{3}=0|yx_{1}=00))$, and $\theta_{7}=f(P(x_{3}=0|yx_{1}=01))$, respectively. While the rotation angles of $C^{2}R_{y}$ gates acting on $x_{2}$ are set as $\theta_{8}=f(P(x_{2}=0|yx_{3}=01))$, $\theta_{9}=f(P(x_{2}=0|yx_{3}=00))$, $\theta_{10}=f(P(x_{2}=0|yx_{3}=10))$, and $\theta_{11}=f(P(x_{2}=0|yx_{3}=11))$, respectively. The output of the TAN structure SN-QBC is
@@
(b) The circuit of the TAN-QBC.
Figure 5: The quantum circuit of the TAN-QBC based on the TAN structure Bayesian network.
III.4 SN-QBC Based on the Symmetric Relationship of Image Attributes
Both the naïve Bayes classifier and the semi-naïve Bayes classifier only consider the dependencies between a small number of attribute nodes. Note that there are symmetric relationships among the sampled feature attributes in some images, such as the digits “6” and “9” shown in Fig. 6, where there exists symmetric relationships between attributes $x_{2}$ and $x_{5}$, and $x_{3}$ and $x_{4}$. These symmetric relationships of image attributes can be used to build Bayesian networks, which can then be used to establish corresponding SN-QBCs. The method for constructing a Bayesian network based on the symmetric relationships of image attributes is as follows.
  1. (1)
@@ III.4 SN-QBC Based on the Symmetric Relationship of Image Attributes
  2. (2)
    Consider the symmetric relationship of attributes in the sample images and add a directed edge between each pair of attributes that are symmetrical to each other.
@@
For the Bayesian network with symmetric relationships between attributes, in the case of not considering the label node $y$, the network will consist of several independent trees. In this case, one can construct the corresponding symmetric-QBC by using the method for the TAN-QBC in Sect. III.3. That is,
  1. (a)
    Each node is initialized to |$\ket{0}$.
  2. (b)
    The class-prior probability $P(y=0)$ is encoded by $R_{y}(\cdot)$.
  3. (c)
    Each independent tree is considered separately. For each tree, starting from the root node, the class-conditional probabilities of each node are encoded by $CR_{y}(\cdot)$ layer by layer, respectively.
Figure 6: The Bayesian network based on symmetric relationships of feature’s attributes.
@@ IV Image Classification Based on QBCs
The key to the accuracy of a Bayesian classifier lies in feature selection and the structure of Bayesian networks. In this paper, we propose an image classification framework based on QBCs and local feature sampling, as illustrated in Fig. (7). Firstly, some local areas are selected in an image for feature sampling. Then, the sampled pixels are pooled and binarized to obtain local binary attributes of features. Finally, a QBC can be constructed based on a Bayesian network model with local attributes as nodes for image classification.
@@
IV.1 Local Feature Sampling
For image classifications using a naïve QBC, if each pixel of the image is taken as a node of a Bayesian network, the network will be very complex. For example, a 28*28 image requires 784 nodes to represent the network. Although $n$ qubits can form a feature space with a dimension of $2^{n}$, the resources required for the quantum Bayesian networks are still relatively large. In addition, considering the complex relationship of interdependence among feature’s attributes, the complexity of Bayesian networks will continue to increase, which hinders the implementation of QBCs on current NISQ devices.
To reduce computational complexity and utilize features effectively for classification, we propose the local feature sampling method. This method aims to obtain a small number of local key attributes from an image. For image classification, background pixels shared by some images might not provide useful information for accurate classification by Bayes classifiers. The local feature sampling method can reduce the number of attribute nodes in Bayesian networks, decrease the scale and computational complexity of QBCs, and diminish the influence of quantum noise, which is beneficial for the experimental implementation of quantum circuits.
@@ IV.1 Local Feature Sampling
  1. (1)
    Local sampling and average pooling. Using a convolution operation on the sampling block and the convolution kernel, an image is locally sampled according to the specified block size (convolution kernel size). Next, the appropriate average pooling value $x^{\prime}_{i}$ of the attribute node $x_{i}$ in the Bayesian network is obtained by applying average pooling YuWang-424 .
  2. (2)
    Feature binarization. In a classical Bayes classifier, assume that the probability density function $p(x_{i}|y)\sim N(\mu_{y,i},\sigma_{y,i}^{2})$, where $\mu_{y,i}$ and $\sigma_{y,i}^{2}$ are the mean and variance of the attribute $x_{i}$ on the $y$-th class. One can obtain the class-prior probability by calculating the probability density function. However, it is challenging to replicate this process for a QBC. In this case, binarization is used to transform $x_{i}$ into 0 or 1 such that it can be represented by a single-qubit.
    Here, we use the maximum likelihood estimation (MLE) Dutilleul-453 ; BoullE-425 method adopted in classical Bayes classifiers for obtaining class-conditional probabilities of continuous variables. For the binary classification problem, this method runs as follows. Assume that $p(x_{i}|y_{0})\sim N(\mu_{y_{0},i},\sigma_{y_{0},i}^{2})$ and $p(x_{i}|y_{1})\sim N(\mu_{y_{1},i},\sigma_{y_{1},i}^{2})$. When two Gaussian functions have only one intersection, which is denoted as $x_{\text{ins}}$, as shown in Fig. (8a), the attribute value is set as follows
    $x_{i}=\begin{cases}0,&\text{if}\ x^{\prime}_{i}\leq x_{\text{ins}}\ \text{and}\ \mu_{y_{0},i}\leq\mu_{y_{1},i},\ \text{or}\ x^{\prime}_{i}>x_{\text{ins}}\ \text{and}\ \mu_{y_{0},i}>\mu_{y_{1},i};\\ 1,&\text{otherwise.}\end{cases}$ (19)
    where $x^{\prime}_{i}$ is the average pooling value generated by the the average pooling processing. If there are two intersections $x_{\text{ins1}}$ and $x_{\text{ins2}}$ with $x_{\text{ins1}}\leq x_{\text{ins2}}$, as shown in Fig. (8b), in cases where $\mu_{y_{0},i}\leq\mu_{y_{1},i}$, the attribute value is set as
@@
    Conversely, when $\mu_{y_{0},i}>\mu_{y_{1},i}$, the attribute value is set as
@@
By utilizing the aforementioned method, it is possible to obtain a limited number of local attributes of features. These local attributes serve as nodes of Bayesian networks for the constructions of QBCs.
IV.2 Image Classification Algorithm Based on QBCs
The image classification algorithm based on QBCs and local feature sampling is shown in Fig. (7). The algorithm consists of two stages, namely the image preprocessing stage, and the Bayesian networks and QBCs construction stage. In the image preprocessing stage, certain local areas are chosen for sampling. After that, the sampled attributes are pooled, and the Gaussian method is used for binarization. This process converts the sampled key attributes into either 0 or 1, allowing an attribute to be represented by a single-qubit in QBCs. In the second stage, a Bayesian network model is selected and the corresponding QBC is constructed with local key attributes as nodes. The algorithm runs as follows.
  1. (a)
@@ IV.2 Image Classification Algorithm Based on QBCs
  2. (b)
    Local sampling and average pooling are performed on images from the training set, and the mean value $\mu_{i}$ and variance $\sigma_{i}$ of the corresponding attributes are calculated. The Gaussian binarization method is executed to obtain the binarized attributes, as described in Sect. IV.1.
  3. (c)
    The QBC circuit is constructed based on the chosen Bayesian network model by using local binarized attributes as nodes. Also see Sect. III.
  4. (d)
    The class-prior probability and class-conditional probability required for the Bayesian network are calculated statistically. These values are then loaded into the QBC circuit as controlled angles to complete the construction of the QBC.
  5. (e)
    To predict the class of a new image, one needs to repeat the step b to obtain a new local key feature $X^{\prime}$ of the image, and then measures the probabilities of states |$\ket{0X^{\prime}}$ and |$\ket{1X^{\prime}}$ on the QBC circuit. The class with the highest probability will be the classification result of the image.
@@ V Simulations
In this paper, we simulate four QBCs on the MindQuantum Mindquantum-418 quantum simulation platform. And the performances of these QBCs are tested using the MNIST LecunBottou-417 and Fashion-MNIST XiaoRasul-416 datasets. In the preprocessing stage, the sampling block size is set to 7*7, and the average pooling method is applied. All images are sampled according to the sampling positions shown in Fig. (9) to obtain 9 local key attributes ($n=9$).
As is shown in Fig. (9), four Bayesian network models are used to construct four corresponding QBCs, i.e., the naïve QBC, the SN-QBC based on the SPODE structure (SPODE-QBC), the SN-QBC based on the TAN structure (TAN-QBC), and the SN-QBC based on the symmetric relationships of attributes (symmetric-QBC). For the TAN structures of the MNIST and Fashion-MNIST datasets, the spanning trees of the training data of the two datasets are calculated according to the maximum weighted spanning tree algorithm Camerini-452 ; Zhou-423 introduced in Sect. III.3.